Lane Company manufactures a single product that requires a great deal of hand labor. Overhead cost is applied on the basis ofCompute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements. (Round your answePrepare a standard cost card for the companys product. (Round your answers to 2 decimal places.) Direct materials pounds atCompute the standard direct labor-hours allowed for the years production. Standard direct labor hoursComplete the following Manufacturing Overhead T-account for the year. Manufacturing OverheadDetermine the reason for any underapplied or overapplied overhead for the year by computing the variable overhead rate and ef
Answer #1

Part 1

Predetermined overhead rate

$15.20

Variable rate

$5.60

Fixed rate

$9.60

Fixed rate = 2880000/300000 = 9.60

Part 2

Direct materials

4

Pounds at

$12

Per pound

$48.00

Direct labor

1.5

DLHs at

$13.80

Per DLH

20.70

Variable overhead

1.5

DLHs at

$5.60

Per DLH

8.40

Fixed overhead

1.5

DLHs at

$9.60

Per DLH

14.40

$91.50

Part 3 a

Standard direct labor hours

360000

240000*1.5 = 360000

Part 3 b

Manufacturing overhead

Actual costs

4368000

5472000

Applied costs

1104000

Overapplied overhead

1248000+3120000 = 4368000

Applied costs = 360000*15.20 = 5472000

Part 4

Variable overhead rate variance

$936000

F

Variable overhead efficiency variance

$168000

U

Fixed overhead budget variance

$240000

U

Fixed overhead volume variance

$576000

F

Variable overhead rate variance = actual overhead –(AH*SR) = 1248000-(390000*5.60)) = -936000

Variable overhead efficiency variance = SR*(AH-SH) = 5.60*(390000-360000) = 168000

Fixed overhead budget variance = actual overhead – budgeted overhead = 3120000-(300000*9.60) = 240000

Fixed overhead volume variance = (300000*9.60)-(360000*9.60) = -576000